Circular No. 7839

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bj IN NGC 1821 Independent discoveries of a supernova in NGC 1821 have been re-

ported by T. Puckett, Mountain Town, GA, with J. Newton, Chiefland, FL, on unfiltered CCD frames taken with the Puckett Observatory 0.60-m automated patrol telescope on Feb. 28.06 (mag 14.8) and Mar. 1.05 UT, and by M. Papenkova and W. D. Li, University of California at Berkeley, on unfiltered CCD LOTOSS (cf. IAUC 7514) images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Feb. 28.2 and Mar. 1.2 (mag ∼ 14.7 on both dates). Puckett measures the new object at α = 5^{h}11^{m}46^{s}.41, δ = −15^{o}08′10′′.8 (equinox 2000.0), which is 4′′.5 east and 5′′.8 south of the center of NGC 1821; Li provides position end figures 46^{s}.37, 10′′.7. Puckett notes that SN 2002bj is not present on Palomar Sky Survey images taken on 1990 Nov. 25 (limiting mag ∼ 21.0) and 1983 Nov. 6 (limiting mag ∼ 20.3), and Li remarks that nothing was visible at this location on a KAIT image taken on 2002 Feb. 21.2 (limiting mag ∼ 18.0).

SUNGRAZING COMETS Further to IAUCs 7797 and 7807, D. Biesecker and D. Hammer re-

port their measurements for additional Kreutz sungrazing comets found on SOHO website images by R. Kracht (the 1999 comets), M. Meyer (C/2002 C3), T. Hoffman (C/2002 C4), and X. Leprette (C/2002 D1). The 2002 comets were detected only with the C3 coronagraph; the 1999 comets were detected only in C2 images. C/2002 C3 showed a tail at apparent distance 25R_{⊙}, while C/2002 C4 was obscured at < 9.5R_{⊙} by a coronal mass ejection. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  1999 UT  α_{2000}  δ_{2000}  MPEC
C/1999 H8 Apr. 19.938  1^{h}55.^{m}5  +10^{o}19′  2002-D42
C/1999 H9  26.254  2 15.7  +11 54  2002-D42
C/1999 J13 May 13.993  3 24.5  +17 03  2002-D45
Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 C3 Feb. 4.654  21^{h}41.^{m}4  −19^{o}00′  2002-C28
C/2002 C4  10.906  21 56.6  −15 36  2002-C95
C/2002 D1  22.029  22 40.9  −11 33  2002-D45
2002 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green